Help! The sum of two integers is 10. Three times the larger integer is three less than eight times the smaller integer. Find the integers.
Answer by kiru_khandelwal(79)

Let the two integers be x and y
Sum of two integers = 10
=> x+y = 10.......eq(1)
Let x > y
So , three time the larger integer is three less then eight times the smaller integer
=> 3x = 8y -3........eq(2)
From equation(1)
x = 10-y
substituting the value of x in eq(2)
3(10-y) = 8y -3
=> 30 -3y = 8y -3
=> 33 = 11y
=> y = 3
substituting the value of y in eq(1) to get value of x
=> x = 10-3 =7
So the two integers are 7 and 3